OpenStudy (anonymous):

Which of these numbers is between 1 and 3/5 8/5 5/9 5/8 3/8

OpenStudy (anonymous):

(3/5) = 0.6 8/5 = 1.6 5/9 = 0.55555555 5/8 = 0.625 3/8 = 0.375

OpenStudy (anonymous):

So, 5/8